# AutoV: Learning to Retrieve Visual Prompt for Large Vision-Language Models

### v1

@@ 4.3.4 The Efficiency of AutoV
While we enhance LVLM performance through AutoV, we uphold efficiency optimizations to ensure practicality. Initially, training a high-quality AutoV on the LLaVA-7B model within epochs only requires hours utilizing NVIDIA A-GB GPUs. Furthermore, because of its transferability, training additional AutoV instances for each vision-language model is unnecessary, and we can generate the training samples via smaller models. Lastly, although AutoV introduces minimal additional FLOPs during inference (e.g., multiple image pre-processing and visual feature extraction), we can reduce latency by downsizing the final similarity matrix and single-prompt selection before LLM processing. As a result, the inference latency of the LLaVA-7B model increases by just 6.4 ms (from 254.8 ms to 261.2 ms), while yielding a 1.3% gain in average accuracy.
@@ 4.3.5 Qualitative Visualization
To further validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses where the black mask removes the background. While referring to the "intended effect" in the image, AutoV selects , which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix G.2.
@@ 5 Conclusion
In this paper, we introduced AutoV, an adaptive visual prompting framework designed to effectively retrieve optimal visual prompts tailored to textual queries and image contexts for large vision-language models (LVLMs). By leveraging a lightweight ranking network trained via a reward-based supervision strategy, AutoV significantly surpasses existing heuristic prompting methods. Extensive experiments across multiple LVLM architectures and benchmarks demonstrate consistent improvements. Future research could explore integrating AutoV with more advanced prompt generation techniques or extending its framework to broader multimodal contexts beyond vision-language scenarios.
Appendix
Appendix A Attention Visual Prompts used in Experiments
@@
Alignment in Vision Foundation Models. As vision foundation models advance rapidly [40, 35, 59, 21], pretrained on language-image pairs, they serve as the visual encoders for large vision-language models. A representative model is CLIP [40], , which consists of a vision encoder and a text encoder, mapping the image and text prompt into the aligned latent space. Therefore, we can directly obtain the similarity between vision modality and text modality feature with .
@@
Attention Visual Prompts. Inspired by the idea that text-image attention (or similarity ) can be visualized as visual prompts for LVLMs [55, 30], where similarity heatmaps overlaid on the image highlight key regions requiring greater attention for prompt-based reasoning. In this work, we adopt a fine-grained similarity visualization manner in API [55], with the -th layer similarity defined as follows:
@@
where denotes the number of transformer layers, is the projection layer with a fully-connected layer and normalization operation, and MSA stands for multi-head self-attention structure. is input vision tokens for the -th layer, while indicates the class token in the token sequence .
Appendix B Training Data
@@
Our training source dataset is composed of several components used in LLaVA [31] to ensure that AutoV is trained on a diverse range of image-question pairs. It includes K samples from the COCO train dataset [28], K samples from the Visual Genome (VG) dataset [23], K samples from the GQA dataset [19], K samples from the OCRVQA dataset [34], and K samples from the TextVQA dataset [44]. In total, the source data contains K image-question pairs. Here, we visualize the proportional relationship of the source dataset in Figure 6 (a). Furthermore, the question types in our training dataset in visualized in Figure 6 (b), and the selected data types are relatively balanced, which enhances the generalizability and robustness of the ranking network in AutoV.
@@
We then generate various visual prompts based on the above source data. Here, we simply use the attention prompting method employed by the API [55] for high-quality generation. To streamline the process, we empirically generate heat maps of attention maps corresponding to CLIP [40] layers , , , and under its setting on the image as our training visual prompts. A specific example (including JSON dictionary and images) can be found in Figure 7. As a result, one question will correspond to four visual prompts, with a total of AutoV being trained on K visual prompts.
Appendix C Evaluation Benchmarks
We conducted experiments on several widely used visual understanding benchmarks.
@@ Appendix C Evaluation Benchmarks
MMMU. [58] The MMMU is designed to evaluate large vision-language models across a wide range of academic disciplines and complex tasks. It covers over 30 subjects including history, medicine, physics, and more, requiring multimodal comprehension, fine-grained reasoning, and domain-specific knowledge. Each question in MMMU is paired with an image and a multiple-choice question, testing the model’s ability to integrate visual understanding with deep textual reasoning.
@@
GQA. [19] The GQA benchmark is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
@@
POPE. [27] The POPE benchmark is primarily used to evaluate the degree of Object Hallucination in models. It reformulates hallucination evaluation by requiring the model to answer a series of specific binary questions regarding the presence of objects in images. Accuracy, Recall, Precision, and F1 Score are effectively employed as reliable evaluation metrics to precisely measure the model’s hallucination level under three different sampling strategies.
@@
TextVQA. [44] The TextVQA benchmark focuses on the comprehensive integration of diverse text information within images. It meticulously evaluates the model’s text understanding and reasoning abilities through a series of visual question-answering tasks with rich textual information. Models need to not only understand the visual content of the images but also be able to read and reason about the text within the images to answer the questions accurately.
@@
MMVet. [57] The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model being able to integrate different core vision-language capabilities. MM-Vet defines 6 core VL capabilities and examines the 16 integrations of interest derived from the capability combination.
@@
LLaVA-In-The-Wild. [31] The LLaVA-In-The-Wild benchmark is built to assess the generalization ability of vision-language models in real-world, unconstrained settings. Unlike curated benchmarks, it features user-collected image-question pairs from diverse sources and domains. This setup presents models with noisy, complex, and unpredictable visual content, offering a realistic measure of their robustness, grounding capability, and performance in open-ended scenarios.
Appendix D Implementation Details.
@@
(1) Data generation stage: LLaVA-v1.5 is consistently adopted as the reference LVLM for generating loss due to resource limitations. (2) AutoV training stage: LLaVA series are trained using DeepSpeed [42] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL and InternVL2 directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
Appendix E Limitations
The introduced AutoV offers a new perspective on retrieving various visual prompts for LVLMs when faced with different queries. However, due to resource constraints, we only generate the training data via LLaVA-7B model [31], and AutoV might yield superior results with more powerful large vision-language models. Besides, to maintain parameter efficiency, we design our approach as a plug-in module for existing LVLMs, requiring no fine-tuning of the model. We note, however, that naively reusing the LLM frozen parameters may lead to suboptimal performance. Finally, due to current limitations in visual prompt quality, our selection is constrained to attention visual prompts [55], resulting in bounded accuracy gains. We believe that when future advances in visual prompt generation, AutoV can further boost their integrated performance.
Appendix F Broader Impacts
Overall, this research has broader impacts on the visual prompt, feature retrieval, and future development of LVLMs, fostering progress and advancements in the field of vision-language models.
Advancing Method: We propose the AutoV model that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image.
Novel Insights: We are the first to apply the reward modeling loss from RLHF [41] to optimal visual prompt retrieval and validate its effectiveness compared with the Mixture-of-Experts framework.
Inspiring Future Research: This work advances visual prompt retrieval in LVLMs, establishing a foundation for future research. By framing the prompt selection challenge systematically, we enable subsequent exploration of (1) higher-quality prompt generation techniques and (2) more effective retrieval methodologies, collectively pushing the boundaries of multimodal understanding.
Appendix G More Visualization
G.1 Cases of Training Data
Visualization of a case in our training data is in Figure 7, including a JSON dictionary and images.
G.2 Retrieval Strategy of AutoV
Visualization of the retrieval strategy employed by AutoV is in Figure 5, with a total of ten cases.

### v2

@@ 4.3.5 The Efficiency of AutoV
While we enhance LVLM performance through AutoV, we uphold efficiency optimizations to ensure practicality. Initially, training a high-quality AutoV on the LLaVA-7B model within epochs only requires hours utilizing NVIDIA A-GB GPUs. Furthermore, because of its transferability, training additional AutoV instances for each vision-language model is unnecessary, and we can generate the training samples via smaller models. Lastly, although AutoV introduces additional TFLOPs during inference (e.g., multiple image pre-processing and visual feature extraction), we can reduce latency by downsizing the final similarity matrix and single-prompt selection before LLM processing. As a result, the inference latency of the LLaVA-7B model increases by just ms (from ms to ms), while yielding a gain in average accuracy.
@@ 4.3.6 Qualitative Visualization
To validate the effectiveness of our retrieval strategy, we visualize an example sourced from LLaVA-Bench within {API-L-Dark, RedCircle, API-L-White, FGVP}. As shown in Figure 5, we list the output results of AutoV when we enter different queries. When faced with a question that focuses on the global context information, AutoV chooses (API-L-Dark), where the black mask removes the background. While referring to the ”intended effect” in the image, AutoV selects (RedCircle), which directly uses red circles to highlight the weird spots. More visualization retrieval cases can be found in the Appendix F.
@@ 5 Conclusion
In this paper, we introduce AutoV, an adaptive visual prompting framework that retrieves optimal visual prompts tailored to textual queries and image contexts for large vision-language models. By leveraging a lightweight ranking network trained with reward-based supervision, AutoV consistently outperforms existing heuristic prompting methods. Extensive experiments across multiple LVLM architectures and benchmarks demonstrate its effectiveness and generality. Overall, AutoV provides a practical and robust solution for enhancing LVLMs’ performance through adaptive visual prompting. Future work explores integrating AutoV with more advanced prompt generation techniques or extending its framework to broader multimodal settings beyond vision-language tasks.
Supplementary Material
Contents
A Details of Training Data
| Method | CVBench | BLINK | RealWorldQA | MMVP | COCOCap | RefCOCO+ | ImageNet-1K | Avg. |
| LLaVA-OneVision (7B) | ||||||||
| Base Model | 79.6 | 48.2 | 66.0 | 60.7 | 61.1 | 68.8 | 92.3 | |
| API (ECCV24) | 80.7 | 49.4 | 66.4 | 61.5 | 61.9 | 69.4 | 92.6 | |
| AutoV | 83.0 | 51.6 | 68.9 | 64.2 | 63.7 | 71.8 | 95.0 | |
| Qwen2.5-VL (7B) | ||||||||
| Base Model | 80.6 | 56.4 | 68.5 | 61.3 | 40.7 | 72.6 | 94.7 | |
| API (ECCV24) | 81.1 | 57.3 | 69.2 | 61.7 | 41.4 | 73.5 | 95.4 | |
| AutoV | 83.0 | 59.0 | 71.1 | 64.3 | 43.3 | 75.2 | 97.1 | |
A.1 Attention Visual Prompts for Training
@@
As vision foundation models advance rapidly [CLIP, oquab2023dinov2, siglip, kirillov2023segment], pretrained on language-image pairs, they serve as the visual encoders for large vision-language models. A representative model is CLIP [CLIP], , which consists of a vision encoder and a text encoder, mapping the image and text prompt into the aligned latent space. Therefore, we can directly obtain the similarity between vision modality and text modality feature with .
@@
Inspired by the idea that text-image attention (or similarity ) can be visualized as visual prompts for LVLMs [api, lin2024training], where similarity heatmaps overlaid on the image highlight key regions requiring greater attention for prompt-based reasoning. In this work, we adopt a fine-grained similarity visualization manner in API [api], with the -th layer similarity defined as follows:
@@
where denotes the number of transformer layers, is the projection layer with a fully-connected layer and normalization operation, and MSA stands for multi-head self-attention structure. is input vision tokens for the -th layer, while indicates the class token in sequence .
A.2 The Composition of Training Data
@@
Our training source dataset is composed of several components used in LLaVA [liu2023improvedllava] to ensure that AutoV is trained on a diverse range of image-question pairs. It includes K samples from the COCO train dataset [lin2014microsoft], K samples from the Visual Genome (VG) dataset [krishna2017visual], K samples from the GQA dataset [hudson2019gqa], K samples from the OCRVQA dataset [mishra2019ocr], and K samples from the TextVQA dataset [singh2019towards]. In total, the source data contains K image-question pairs. Here, we visualize the proportional relationship of the source dataset in Figure 6 (a). Furthermore, the question types in our training dataset in visualized in Figure 6 (b), and the selected data types are relatively balanced, which enhances the generalizability and robustness of the ranking network in AutoV.
@@
We then generate various visual prompts based on the above source data. Here, we simply use the attention prompting method employed by the API [api] for high-quality generation. To streamline the process, we empirically generate heat maps of attention maps corresponding to CLIP [CLIP] layers , , , and under its setting on the image as our training visual prompts.
A specific example (including JSON dictionary and images) can be found in Figure 7. As a result, one question will correspond to four visual prompts, with a total of AutoV being trained on K visual prompts.
B Evaluation Benchmarks
We conduct comprehensive experiments on fifteen widely used vision-language benchmarks.
@@ B Evaluation Benchmarks
MMMU. [yue2024mmmu] The MMMU is designed to evaluate LVLMs across a wide range of academic disciplines and complex tasks. It covers over subjects, including history, medicine, physics, and more, requiring multimodal comprehension, fine-grained reasoning, and domain-specific knowledge. Each question is paired with an image and a multiple-choice question, testing the ability to integrate visual understanding with deep textual reasoning.
@@
GQA. [hudson2019gqa] The GQA benchmark is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
@@
POPE. [li2023evaluating] The POPE benchmark is primarily used to evaluate the degree of Object Hallucination in models. It reformulates hallucination evaluation by requiring the model to answer a series of specific binary questions regarding the presence of objects in images. Accuracy, Recall, Precision, and F1 Score are effectively employed as reliable evaluation metrics to precisely measure the model’s hallucination level under three different sampling strategies.
@@
TextVQA. [singh2019towards] The TextVQA benchmark focuses on the comprehensive integration of diverse text information within images. It meticulously evaluates the model’s text understanding and reasoning abilities through a series of visual question-answering tasks with rich textual information. Models need to not only understand the visual content of the images but also be able to read and reason about the text within the images to answer the questions accurately.
@@
MMVet. [yu2023mm] The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model that can integrate different core vision-language capabilities. MM-Vet defines 6 core capabilities and examines the 16 integrations of interest derived from the capability combination.
@@
LLaVA-In-The-Wild. [liu2023improvedllava] The LLaVA-In-The-Wild benchmark is to assess the generalization ability of LVLMs in real-world, unconstrained settings. Unlike curated benchmarks, it features user-collected image-question pairs from diverse sources and domains. This setup presents models with noisy, complex, and unpredictable visual content, offering a realistic measure of their robustness, grounding, and performance in open-ended scenarios.
MMStar. [chen2024we] MMStar evaluates six vision-language capabilities across 18 fine-grained dimensions using visually dependent questions. It introduces Multi-modal Gain and Multi-modal Leakage metrics to measure visual contribution and data leakage. This benchmark rigorously tests models’ true visual reasoning and generalization abilities.
MathVista. [lu2023mathvista] MathVista assesses mathematical reasoning grounded in visual contexts such as diagrams and plots. It integrates 28 prior datasets and adds new subsets targeting logical, functional, and scientific figure reasoning. The benchmark challenges models to combine precise visual understanding with compositional calculations and multi-step quantitative reasoning.
CVBench. [cvbench2025] CVBench measures models’ ability to reason across multiple videos. Its tasks involve object association, event association, and complex multi-video reasoning between distinct video clips. The dataset highlights large performance gaps between humans and models in multi-video reasoning and temporal integration.
BLINK. [fu2024blink] BLINK targets core perceptual abilities that humans solve instantly but multimodal models struggle with. It reformulates 14 classic vision tasks into 3,800 multiple-choice questions testing fine-grained perception. Results reveal that current models still lack fundamental visual understanding compared to humans.
Real-World QA. [realworldqa2025] In order to develop useful real-world AI assistants, it is crucial to advance a model’s understanding of the physical world. RealWorldQA is a benchmark just for the design. The dataset consists of anonymized images taken from vehicles, in addition to other real-world images. The initial release of the RealWorldQA consists of over 700 images, with a question and easily verifiable answer for each image.
MMVP. [tong2024eyes] MMVP exposes perceptual blind spots in vision-language models using “CLIP-blind” image pairs. These are pairs of images that contain clear visual differences to humans but are nearly indistinguishable to CLIP-based vision encoders. Leveraging such tricky image pairs, the benchmark pinpoints core weaknesses in visual representation and perception fidelity.
COCOCap. [chen2015microsoftcococaptionsdata] Microsoft COCO Image Captioning benchmark, a standard test of a model’s ability to generate natural language descriptions of images. It contains over 330,000 images with five human-written captions each, scored by metrics like BLEU [papineni-etal-2002-bleu] and CIDEr [vedantam2015ciderconsensusbasedimagedescription]. Performance reflects a model’s capacity for coherent and contextually accurate visual description.
RefCOCO+. [kazemzadeh-etal-2014-referitgame] RefCOCO+ measures models’ ability to ground appearance-based referring expressions in images. Its annotations forbid spatial words, forcing reliance on attributes. This benchmark tests precise visual grounding under complex, attribute-driven conditions.
ImageNet-1K. [imagenet15russakovsky] ImageNet-1K is a foundational benchmark for large-scale image classification. It contains 1,000 object categories with 1.28M training and 50K validation images. Top-1 and Top-5 accuracy serve as key metrics for evaluating visual recognition performance.
C Implementation Details
@@
(1) Data generation stage: LLaVA-v1.5 is consistently adopted as the reference LVLM for generating loss due to resource limitations. (2) AutoV training stage: The LLaVA series is trained using DeepSpeed [rasley2020deepspeed] ZeRO2 with A100-40G GPUs within epochs. The batch size (with accumulation) is set to and the learning rate is -. For convenience, Qwen2.5-VL, InternVL2, Gemini-1.5-Pro, GPT-4o, and Qwen-VL-Max directly adopt the retrieval strategy in AutoV pre-trained from LLaVA-OneVision.
D More Main Experiments
In this section, we present the extended results that complement the main experiments reported in Section 4. Specifically, we provide detailed quantitative results and additional analyses of AutoV on a broader range of benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classical tasks such as COCO2017 Captioning, RefCOCO+ Grounding, and ImageNet-1K Classification. All experiments are conducted using LLaVA-OneVision 7B and Qwen2.5-VL 7B as base models. The results consistently demonstrate that AutoV effectively enhances multimodal understanding and reasoning across diverse domains, confirming its model-agnostic generalization capability.
| Run | Seed | AutoV | Baseline | Diff. |
| 1 | 10 | 38.7 | 36.3 | +2.4 |
| 2 | 20 | 38.7 | 36.3 | +2.4 |
| 3 | 30 | 38.8 | 36.2 | +2.6 |
| 4 | 40 | 38.7 | 36.2 | +2.5 |
| 5 | 50 | 38.8 | 36.4 | +2.4 |
E Statistical Significance Analysis
To further verify the robustness of the improvement in the main experiments, we conduct significance testing across multiple random seeds on representative benchmarks. Specifically, we present the detailed analysis process for MMMU as an example, while the same manner applies to the remaining benchmarks. The results are shown in Table 6, and the analysis is presented below:
Mean : +2.06%
Std Dev:
-value:
The above analysis provides solid statistical evidence that AutoV delivers consistent and reproducible improvements across runs, highlighting its robustness and reliability in enhancing vision-language model performance.
F More Visualization Results
F.1 Cases of Training Data
Figure 7 shows one training example in our dataset, including its JSON structure and the corresponding attention maps. The JSON entry stores the image path, human–GPT conversation, different visual prompt paths from the API, as well as the predicted ranking and loss values. The right side visualizes these visual prompts, illustrating how different prompts emphasize varying regions of the same image.
F.2 Retrieval Strategy of AutoV
To further demonstrate the effectiveness of our retrieval strategy, Figure 8 provides additional visualizations of ten representative cases beyond those presented in the main text, showing how AutoV retrieves the most suitable visual prompt according to different textual queries. These examples are sourced from different benchmarks, covering multiple visual prompt sets, including {API-L-Dark, RedCircle, API-L-White, FGVP}. Best viewed in color.